Compare the fractions and put an appropriate sign.
 1 / 7 ☐ 1 / 4

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Multiply by 4
1 / 7 = (1 × 4) / (7 × 4)
= 4 / 28
Multiply by 7
1 / 4 = (1 × 7) / (4 × 7)
]= 7 / 28
Here 4 < 7
∴1 / 7 < 1 / 4

​Thanks